OpenStudy (anonymous):

If f(x) = 5x + 3, then what is f-1(x)

jimthompson5910 (jim_thompson5910):

f(x) = 5x + 3 y = 5x + 3 x = 5y + 3 Now solve for y

OpenStudy (anonymous):

I dont understand

jimthompson5910 (jim_thompson5910):

What do you get when you solve for y

jimthompson5910 (jim_thompson5910):

I started off by replacing f(x) with y. Then I swapped x and y. The goal now is to solve for y.

OpenStudy (anonymous):

plug in what you got for x

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

plug in what I got for x into the y= equation

jimthompson5910 (jim_thompson5910):

no, just tell me what you get when you solve for y

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

solve x = 5y + 3 for y

OpenStudy (anonymous):

y = 3/-5

jimthompson5910 (jim_thompson5910):

\[\Large x = 5y + 3\] \[\Large x-3 = 5y\] \[\Large \frac{x-3}{5} = y\] \[\Large y=\frac{x-3}{5} \] So the inverse function is \[\Large f^{-1}(x)=\frac{x-3}{5} \]

OpenStudy (anonymous):

thank you but where did the f^-1 come from

jimthompson5910 (jim_thompson5910):

that stands for the inverse function

jimthompson5910 (jim_thompson5910):

once you've solved for y, that's effectively finding the inverse

OpenStudy (anonymous):

ok. Thanks for the help

jimthompson5910 (jim_thompson5910):

np